The healthcare provider writes an order for a loading dose of theophylline 5 mg/kg PO daily. The client weighs 100 kg. The medication is supplied 150 mg/15 mL. How many mL would you administer?

A.253 mL

B.21.7 mL

C.50 mL

D.100 mL
Answer #1

5 mg/kg PO means

1kg = 5 mg

since client weighs 100 kg

so 100 kg = 5 * 100 mg = 500 mg

Required dose in mg = 500 mg

medication is supplied 150 mg/15 mL

means

150 mg = 15 mL

1 mg = 15 /150 mL

1 mg = 0.1 mL

so 500 mg = 500 * 0.1 mL = 50 mL

50 mL would you administer